# TiM4Rec: An Efficient Sequential Recommendation Model Based on Time-Aware Structured State Space Duality Model

### v2

@@ 3.4. Time-aware Masked Matrix
$\hat{\bm{A}}_{i}=\hat{a}_{i}=\mathcal{D}_{i}\cdot\bm{A}_{i}=\hat{d}_{ii-1}\cdot a_{i}=d_{i}\cdot a_{i}$
$\hat{\bm{A}}_{m:n}\coloneqq\hat{a}_{n}\hat{a}_{n-1}\cdots\hat{a}_{m}=a_{n}d_{n}\cdot a_{n-1}d_{n-1}\cdots a_{m}d_{m}$
(7) $\bm{L}=\begin{bmatrix}1\\ \hat{a}_{1}&1\\ \hat{a}_{2}\hat{a}_{1}&\hat{a}_{2}&1\\ \vdots&\vdots&\ddots&\ddots\\ \hat{a}_{t-1}\ldots\hat{a}_{1}&\hat{a}_{t-1}\ldots\hat{a}_{2}&\cdots&\hat{a}_{t-1}&1\end{bmatrix}$
This can be regarded as an equivalent approximation of $d_{ij}$ as follows:
$\hat{d}_{ij}\approx\hat{d}_{ii-1}\hat{d}_{i-1i-2}\cdots\hat{d}_{j+1j}=d_{i}d_{i-1}\cdots d_{j}$
In the context of SSD kernel computations, multiplication is effectively converted into accumulation operations via the logarithm function. This transformation offers advantageous support for the aforementioned approximation of $\hat{d}_{ij}$.
$\log\left(d_{i}d_{i-1}\cdots d_{j}\right)=\log\left(d_{i}\right)+\log\left(d_{i-1}\right)+\cdots+\log\left(d_{j}\right)$
Through the approximation method mentioned above, we accomplish the time-aware enhancement of the SSD architecture. This is achieved without disrupting the semi-separable property of the masked matrix $\bm{L}$.
@@ 3.5. Time-aware SSD Block
Despite providing a detailed derivation of the SSD in Section 2.3, the actual code implementation of the SSD kernel retains the discretization process. Therefore, further analysis is required. Notably, in Section 2.2, we derive the discretization for the SSM, but it is important to highlight that the actual computational process is approximated in both SSM and SSD implementations. Specifically, for matrix $\overline{\bm{A}}$, the implementation approximates the exponential operation by directly computing $\overline{\bm{A}}\approx\Delta\bm{A}$. Meanwhile, for matrix $\overline{\bm{B}}$, given the Maclaurin series expansion of $e^{x}$ as $1+x+\frac{x^{2}}{2}+\cdots$, the following approximation can be made:
$\overline{\bm{B}}$ $\approx\left(\Delta\bm{A}\right)^{-1}\left(\bm{I}+\Delta\bm{A}-\bm{I}\right)\cdot\Delta\bm{B}$
@@
$=\Delta\bm{B}$
In this approximation, we establish the following definitions: the state transition coefficient $\bm{A}\in\mathbb{R}^{1}$. By inputting $\bm{X}\in\mathbb{R}^{T\times D}$, we derive matrices $\bm{B}\in\mathbb{R}^{T\times N}$ and $\bm{C}\in\mathbb{R}^{T\times N}$, along with the discretization parameter $\Delta\in\mathbb{R}^{T}$:
$\bm{X},\bm{B},\bm{C},\Delta=\bm{X}W+b$
@@
$\bm{X}_{t},\bm{B}_{t},\bm{C}_{t}$ $=\sigma\left[(\bm{X}_{t},\bm{B}_{t},\bm{C}_{t})^{\top}*\omega\right]$
$\text{where}\quad\mathcal{Q}_{t}=\mathcal{P}_{t}*\omega$ $\coloneqq\sum_{m=0}^{K-1}\mathcal{P}_{\max\left(t-m,0\right)}\cdot\omega_{m}$
@@
$\alpha^{\mathcal{D}}=\sigma\left[\sigma\left(\mathcal{D}W_{1}^{\mathcal{D}}+b_{1}^{\mathcal{D}}\right)W_{2}^{\mathcal{D}}+b_{2}^{\mathcal{D}}\right]$
$\mathcal{D}=\alpha^{\mathcal{D}}\cdot\mathcal{D}$
@@
$\mathcal{D}_{t}=\sigma\left(\mathcal{D}_{t}*\omega^{\mathcal{D}}\right)\coloneqq\sigma\left(\sum_{m=0}^{K-1}\mathcal{D}_{\max\left(t-m,0\right)}\cdot\omega^{\mathcal{D}}_{m}\right)$
@@
$\hat{\Delta}=\Delta\cdot\mathcal{D},\quad\overline{\bm{A}}=\hat{\bm{A}}=\hat{\Delta}\cdot\bm{A},\quad\overline{\bm{B}}=\hat{\Delta}\cdot\bm{B}$
Finally, the following equation can be derived to map the input $\bm{X}$ to the output $\bm{Y}\in\mathbb{R}^{T\times D}$($\bm{L}$ refers to eq. 7.):
$\bm{Y}=\bm{L}\circ\bm{C}\overline{\bm{B}}^{\top}\bm{X}=\bm{L}\circ\bm{C}\bm{B}^{\top}\left(\hat{\Delta}^{\top}\bm{X}\right)$
Due to the preservation of the semi-separable matrix property of $\bm{L}$, the matrix computation acceleration method discussed in Section 2.3 remains applicable.
To extract deeper user interest features, multiple layers of Time-aware SSD blocks are stacked. Consequently, after the Time-aware SSD, the features are transformed by adding a Feed Forward Network layer to adapt to the next semantic space. It is noteworthy that to adapt the time difference vector $\mathcal{D}$ to the feature semantic space of the next layer, gate residual (he2016deep) processing is applied to the input time difference vector $\mathcal{D}$ for the subsequent layer.
@@ 3.6. Prediction Layer
By utilizing the Time-aware SSD Block, input $\bm{X}$ can be transformed into a user interest feature sequence $\bm{P}\in\mathbb{R}^{L\times D}$. The last element $p_{t}$ of the feature sequence $\bm{P}$ is then extracted as the interest feature corresponding to the current user interaction sequence.
$\hat{\mathcal{Y}}=Softmax\left(p_{t}\cdot\mathcal{E}\right)$
As stated in Section 2.1, the predicted scores $\hat{\mathcal{Y}}\in\mathbb{R}^{|\mathcal{V}|}$ for all items are obtained by calculating the inner product of $p_{t}$ and the embedding table of all items. However, a notable difference is that the actual score calculation incorporates the Softmax function, even though the Softmax operation does not influence the ranking of the scores.
@@ 3.7. Complexity Analysis
Models based on the Vanilla Transformer, such as SASRec, require a total of $O(L^{2}N)$ floating-point operations (FLOPs). In contrast, thanks to the computational improvements provided by Mamba, models based on SSM and SSD, such as Mamba4Rec and SSD4Rec, only require $O(LN^{2})$ FLOPs. In practical application scenarios of sequential recommendation, including news and product recommendations, it is often the case that $L\gg N$. Thus, sequential recommendation models based on Mamba demonstrate superior advantages. TiSASRec enhances its performance by integrating time-aware capabilities via time difference encoding. Despite the theoretical FLOPS remaining in the order of $O(L^{2}N)$, the practical computational and memory demands exceed those of SASRec by more than four times. This substantial increase is attributed to the inherent computational constraints associated with time difference encoding. Consequently, while TiSASRec demonstrates improved time sensitivity, it necessitates a significantly higher computational overhead and memory allocation compared to SASRec, posing potential challenges in scalability and efficiency. In contrast, TiM4Rec preserves the benefits of the SSD architecture by employing a time difference masked matrix characterized by linear computational complexity. This approach maintains the computational efficiency, demanding only $O(LN^{2})$ FLOPs, and incurs only a modest increase in additional computational overhead. Notably, the generation of the time-aware masked matrix in TiM4Rec is solely based on scalar computations. This characteristic ensures that the additional computational complexity does not escalate in high-dimensional modeling scenarios, thereby preserving the superior matrix operation efficiency inherent to the SSD4Rec architecture.
Table 1. Comparison of computational complexity between representative sequential recommendation methods, where ”M.M.” refers to ”Matrix Multiplication”.
Architecture FLOPs Memory M.M.
SASRec Attention $L^{2}N$ $L^{2}N$
Mamba4Rec SSM $LN^{2}$ $LN^{2}$
SSD4Rec SSD $LN^{2}$ $LN$
TiSASRec Attention $L^{2}N$ $L^{2}N$
TiM4Rec SSD $LN^{2}$ $LN$
@@ 4. Experiments
Table 2. Sequential recommendation performance on the ML-1M, Beauty, and KuaiRand datasets. Here, R@k refers to Recall@K, N@K denotes NDCG@K, and M@K represents MRR@K. The best and second-best results are indicated in bold and underlined, respectively. ”Improv.” signifies the performance improvement of TiM4Rec compared to the second-best results. The asterisk in SSD4Rec* denotes that this is our own replicated version.
Model Type & Model $\rightarrow$ CNN RNN Transformer SSM SSD Improv.
Dataset Metric Caser GRU4Rec SASRec BERT4Rec TiSASRec Mamba4Rec SSD4Rec* TiM4Rec
MovieLens-1M R@10 0.2156 0.2985 0.3060 0.2800 0.3147 0.3253 0.3199 0.3310 1.75%
R@20 0.3083 0.3937 0.4050 0.3853 0.4250 0.4354 0.4240 0.4338 -0.30%
R@50 0.4515 0.5397 0.5455 0.5377 0.5608 0.5707 0.5649 0.5770 1.10%
N@10 0.1157 0.1700 0.1754 0.1584 0.1834 0.1891 0.1841 0.1932 2.17%
N@20 0.1391 0.1941 0.2005 0.1852 0.2106 0.2169 0.2104 0.2194 1.15%
N@50 0.1675 0.2232 0.2285 0.2154 0.2376 0.2436 0.2384 0.2477 1.68%
M@10 0.0853 0.1307 0.1357 0.1214 0.1424 0.1474 0.1425 0.1512 2.58%
M@20 0.0917 0.1373 0.1425 0.1288 0.1498 0.1551 0.1498 0.1584 2.13%
M@50 0.0963 0.1420 0.1471 0.1337 0.1542 0.1593 0.1542 0.1629 2.26%
Amazon Beauty R@10 0.0402 0.0563 0.0851 0.0352 0.0802 0.0838 0.0806 0.0854 0.35%
R@20 0.0618 0.0856 0.1194 0.0563 0.1147 0.1185 0.1146 0.1204 0.84%
R@50 0.1090 0.1376 0.1759 0.0999 0.1734 0.1802 0.1710 0.1800 -0.11%
N@10 0.0199 0.0297 0.0425 0.0172 0.0405 0.0435 0.0423 0.0446 2.53%
N@20 0.0253 0.0371 0.0511 0.0225 0.0491 0.0522 0.0509 0.0533 2.11%
N@50 0.0346 0.0473 0.0623 0.0311 0.0605 0.0644 0.0620 0.0651 1.09%
M@10 0.0138 0.0217 0.0294 0.0117 0.0282 0.0311 0.0306 0.0321 3.22%
M@20 0.0153 0.0237 0.0317 0.0132 0.0306 0.0335 0.0329 0.0345 2.99%
M@50 0.0167 0.0253 0.0335 0.0145 0.0325 0.0354 0.0347 0.0363 2.54%
Kuai Rand R@10 0.0801 0.1020 0.1055 0.0938 0.1057 0.1094 0.1055 0.1109 1.37%
R@20 0.1344 0.1659 0.1704 0.1537 0.1710 0.1768 0.1717 0.1774 0.34%
R@50 0.2561 0.3017 0.3074 0.2873 0.3060 0.3154 0.3088 0.3202 1.52%
N@10 0.0395 0.0564 0.0584 0.0510 0.0590 0.0608 0.5880 0.0611 0.49%
N@20 0.0531 0.0724 0.0747 0.0660 0.0753 0.0777 0.0754 0.0779 0.26%
N@50 0.0770 0.0911 0.1016 0.0923 0.1019 0.1050 0.1024 0.1060 0.95%
M@10 0.0273 0.0428 0.0443 0.0382 0.0450 0.0461 0.0449 0.0463 0.43%
M@20 0.0310 0.0471 0.0487 0.0422 0.0494 0.0508 0.0494 0.0508 -
M@50 0.0347 0.0513 0.0529 0.0464 0.0536 0.0552 0.0536 0.0552 -
@@ 4.1.1. Datasets
  • MovieLens-1M (harper2015movielens): A dataset containing approximately 1 million user ratings for movies, collected from the MovieLens platform.
  • Amazon-Beauty (mcauley2015image): The user review dataset collected in the beauty category on the Amazon platform was compiled up to the year 2014.
  • Kuai Rand (kuairand): Acquired from the recommendation logs of the application Kuaishou, the dataset includes millions of interactions involving items that were randomly displayed.
For each user, we sort their interaction records according to the timestamps, thereby generating an interaction sequence for each individual. And then retain only those users and items that are associated with a minimum of five interaction records, in accordance with the methodology established in prior research (kang2018self). The statistics of these datasets are shown in Table 3.
4.1.2. Baseline
To demonstrate the superiority of our proposed method, we select a set of representative sequential recommendation models to serve as baselines. Notably, even though SSD4Rec (ssd4rec) has not been released as open-source, we implemented a variant termed SSD4Rec*, adhering to the traditional sequential recommendation framework to validate the efficacy of our time-aware approach. This implementation is strictly based on the SSD architecture and does not incorporate features such as variable-length sequences and bidirectional SSD methods proposed in SSD4Rec. The exclusion of these features is intentional, as they are generally regarded as universal enhancement techniques, particularly in the context of handling variable-length sequences.
  • Caser (tang2018personalized): A classical sequential recommendation model based on CNN.
  • GRU4Rec (GRU4Rec): A RNN-based method constructed by Gated Recurrent Units (GRU).
  • SASRec (kang2018self): The first exploration model of applying Transformer in the field of sequential recommendation.
  • BERT4Rec (sun2019bert4rec): A bidirectional attention sequential recommendation model following the BERT (bert) model paradigm.
  • TiSASRec (TiSASRec): A time-aware enhanced sequential recommendation model based on SASRec: Considered as paradigm in time-aware sequential recommendation research.
  • Mamba4Rec (mamba4rec): A pioneering model that explores the application of the Mamba architecture in the domain of sequential recommendation.
  • SSD4Rec (ssd4rec): The pioneering sequential recommendation model leveraging the SSD architecture, exploits its inherent advantages over the SSM architecture for high-dimensional modeling, and integrates variable-length sequence training techniques to enhance model performance.
4.1.3. Evaluation Metrics

### v3

@@ 3.3.2. Time-aware Structured Masked Matrix
$\hat{\bm{A}}_{i}=\hat{a}_{i}=\mathcal{D}_{i}\cdot\bm{A}_{i}=\hat{d}_{ii-1}\cdot a_{i}=d_{i}\cdot a_{i},$
$\hat{\bm{A}}_{m:n}\coloneqq\hat{a}_{n}\hat{a}_{n-1}\cdots\hat{a}_{m}=a_{n}d_{n}\cdot a_{n-1}d_{n-1}\cdots a_{m}d_{m},$
(12) $\bm{L}=\begin{bmatrix}\hat{a}_{0}\\ \hat{a}_{1}&\hat{a}_{0}\\ \hat{a}_{2}\hat{a}_{1}&\hat{a}_{2}&\hat{a}_{0}\\ \vdots&\vdots&\ddots&\ddots\\ \hat{a}_{t-1}\ldots\hat{a}_{1}&\hat{a}_{t-1}\ldots\hat{a}_{2}&\cdots&\hat{a}_{t-1}&\hat{a}_{0}\end{bmatrix},$
We refer to $\bm{L}$ as a Time-aware Structured Masked Matrix. This can be regarded as an equivalent approximation of $\hat{d}_{ij}$ as follows (Notably, the computation in Eq. 12 employs the segment accumulation, consistent with the 1-SS matrix computation in SSD (SSD2024), to maintain efficient computational performance.):
$\hat{d}_{ij}\approx\hat{d}_{ii-1}\hat{d}_{i-1i-2}\cdots\hat{d}_{j+1j}=d_{i}d_{i-1}\cdots d_{j},$
Through the approximation method mentioned above, we accomplish the time-aware enhancement of the SSD architecture. This is achieved without disrupting the 1-SS property of the masked matrix $\bm{L}$.
Figure 4. In comparing our proposed method with the time-aware enhancement approach based on the TiSASRec paradigm, we emphasize the distinctions in how time difference information is transformed into attention score matrices. The time difference matrix is defined in the upper left corner, as outlined in Eq. 11.
By directly integrating the time difference matrix into the masked matrix using the aforementioned method, computational efficiency is maintained. As illustrated in Figure. 4, unlike the time-aware augmentation methods based on the TiSASRec paradigm (TiSASRec2020), our approach circumvents the need to elevate time difference information to high-dimensional processing. Instead, it employs a straightforward scalar-level sequence transformation (which will be detailed shortly) to effectively capture personalized user interaction patterns. This enhanced time difference information is then utilized to construct a Time-aware Structured Masked Matrix $\bm{L}$ that adheres to the property of 1-SS matrix, integrating seamlessly into the SSD kernel. This approach addresses the dual challenges of computational efficiency and integration into the SSD kernel.
@@ 3.4. Time-aware SSD Block
In the following section, we will elaborate on the integration of the Time-aware Structured Masked Matrix into the SSD kernel. Despite providing a detailed derivation of the SSD in Section 2.4, the actual code implementation of the SSD kernel retains the discretization process. Therefore, further analysis is required. Notably, in Section 2.3, we derive the discretization for the SSM, but it is important to highlight that the actual computational process is approximated in both SSM and SSD implementations. Specifically, for matrix $\overline{\bm{A}}$, the implementation approximates the exponential operation by directly computing $\overline{\bm{A}}\approx\Delta\bm{A}$. Meanwhile, for matrix $\overline{\bm{B}}$, given the Maclaurin series expansion of $\exp(x)$ as $1+x+\frac{x^{2}}{2}+\cdots$, the following approximation can be made:
(13) $\overline{\bm{B}}$ $\approx\left(\Delta\bm{A}\right)^{-1}\left(\bm{I}+\Delta\bm{A}-\bm{I}\right)\cdot\Delta\bm{B}$
@@
$=\Delta\bm{B},$
Under this approximation, the original state transition matrix $\bm{A}\in\mathbb{R}^{T}$ in the SSD can be transformed into a scalar coefficient $A\in\mathbb{R}^{1}$, which when multiplied by the discretization parameter $\Delta\in\mathbb{R}^{T}$, yields the discretized state transition matrix $\overline{\bm{A}}\in\mathbb{R}^{T}$. This approximation process optimizes computational efficiency while maintaining maximal accuracy. Given that this approach is empirically established by Gu et al. (SSD2024), we will not delve into further details here.
Building upon the aforementioned consensus, we now proceed to a comprehensive derivation of the Time-aware SSD Block method. By inputting $\bm{X}\in\mathbb{R}^{T\times D}$, we derive matrices $\bm{B}\in\mathbb{R}^{T\times N}$ and $\bm{C}\in\mathbb{R}^{T\times N}$, along with the discretization parameter $\Delta\in\mathbb{R}^{T}$:
(14) $\bm{X},\bm{B},\bm{C},\Delta=\bm{X}W+b,$
@@
(15) $\bm{X}_{t},\bm{B}_{t},\bm{C}_{t}=\sigma\left[(\bm{X}_{t},\bm{B}_{t},\bm{C}_{t})^{\top}*\omega\right],$
$\text{where}\quad\mathcal{Q}_{t}=\mathcal{P}_{t}*\omega\coloneqq\sum_{m=0}^{K-1}\mathcal{P}_{\max\left(t-m,0\right)}\cdot\omega_{m},$
@@
(16) $\alpha^{\mathcal{D}}$ $=\sigma\left[\sigma\left(\mathcal{D}W_{1}^{\mathcal{D}}+b_{1}^{\mathcal{D}}\right)W_{2}^{\mathcal{D}}+b_{2}^{\mathcal{D}}\right],$
$\mathcal{D}$ $=\alpha^{\mathcal{D}}\cdot\mathcal{D},$
@@
(17) $\mathcal{D}_{t}=\sigma\left(\mathcal{D}_{t}*\omega^{\mathcal{D}}\right)\coloneqq\sigma\left(\sum_{m=0}^{K-1}\mathcal{D}_{\max\left(t-m,0\right)}\cdot\omega^{\mathcal{D}}_{m}\right),$
@@
(18) $\hat{\Delta}=\Delta\cdot\mathcal{D},\quad\overline{\bm{A}}=A\cdot\hat{\Delta},\quad\overline{\bm{B}}=\hat{\Delta}\cdot\bm{B},$
Following the methodology outlined in Section 3.3.2, we employ the following transformations to construct the Time-aware Structured Masked Matrix $\bm{L}$:
$\hat{\bm{A}}_{i}=\hat{a}_{i}=\overline{\bm{A}}_{i}=A\cdot\hat{\Delta}_{i}=A\cdot\Delta_{i}\cdot d_{i},$
(19) $\bm{L}=\begin{bmatrix}\hat{a}_{0}\\ \hat{a}_{1}&\hat{a}_{0}\\ \hat{a}_{2}\hat{a}_{1}&\hat{a}_{2}&\hat{a}_{0}\\ \vdots&\vdots&\ddots&\ddots\\ \hat{a}_{t-1}\ldots\hat{a}_{1}&\hat{a}_{t-1}\ldots\hat{a}_{2}&\cdots&\hat{a}_{t-1}&\hat{a}_{0}\end{bmatrix},$
Finally, the following equation can be derived to map the input sequence $\bm{X}$ and $\mathcal{D}$ to the output $\bm{Y}\in\mathbb{R}^{T\times D}$:
(20) $\bm{Y}=TiSSD\left(\bm{X},\mathcal{D}\right)\coloneqq\bm{L}\circ\bm{C}\overline{\bm{B}}^{\top}\bm{X}=\bm{L}\circ\bm{C}\bm{B}^{\top}\left(\hat{\Delta}^{\top}\bm{X}\right),$
Due to the preservation of the 1-SS matrix property of $\bm{L}$, the matrix computation acceleration method discussed in Section 2.4 remains applicable.
Multiple layers of Time-aware SSD blocks are stacked to extract deeper user interest features. Consequently, after the Time-aware SSD, the features are transformed by adding an FFN (Transformer2017) layer to adapt to the next semantic space. It is noteworthy that to adapt the time difference vector $\mathcal{D}$ to the feature semantic space of the next layer, gate residual (ResNet2016) processing is applied to the input time difference vector $\mathcal{D}$ for the subsequent layer.
@@ 3.5. Prediction Layer
By utilizing the Time-aware SSD Layer, input $\bm{X}$ can be transformed into a user interest feature sequence $\bm{P}\in\mathbb{R}^{T\times D}$. The last element $p\in\mathbb{R}^{D}$ of the feature sequence $\bm{P}$ is then extracted as the interest feature corresponding to the current user interaction sequence. Afterward, we compute the dot product between the user interest feature $p$ and the embeddings of all items $\mathbb{E}\in\mathbb{R}^{|\mathcal{V}|\times D}$. This result is then scaled using the Softmax function to obtain the interaction probability scores $\hat{\mathcal{Y}}\left(v_{T+1}|p\right)\in\mathbb{R}^{|\mathcal{V}|}$ for each item from the user.
(21) $\hat{\mathcal{Y}}\left(v_{T+1}|p\right)=Softmax\left(p\cdot\mathbb{E}^{T}\right),$
In alignment with previous work (SASRec2018), we optimize the model parameters using the Cross Entropy loss function. Let us denote the true interaction distribution for user $u_{k}$ as $g^{u_{k}}\in\mathbb{R}^{|\mathcal{V}|}$; then, the following loss function is defined:
(22) $\mathcal{L}=-\sum_{u_{k}}^{|\mathcal{U}|}\sum_{i}^{|\mathcal{V}|}g^{u_{k}}_{i}\ln\left(\hat{\mathcal{Y}}^{u_{k}}_{i}\right)$
@@ 3.6. Complexity Analysis
Models based on the Vanilla Transformer (Transformer2017), such as SASRec (SASRec2018), require a total of $O(L^{2}N)$ floating-point operations (FLOPs). Models based on Linear Recurrent Units (LRU2023), such as LRURec (LRURec2024), require $O(\log(T)N^{2})$ FLOPs. In contrast, thanks to the computational improvements provided by Mamba, models based on SSM (Mamba2023) and SSD (SSD2024), such as Mamba4Rec (Mamba4Rec2024) and SSD4Rec (SSD4Rec2024), only require $O(TN^{2})$ FLOPs. Notably, although the computational complexity of the SSM and SSD architectures is identical, the SSD architecture can efficiently leverage matrix multiplication for computations, resulting in faster actual training times compared to the SSM architecture. In practical application scenarios of sequential recommendation, including news and product recommendations, it is often the case that $T\gg N$. Thus, sequential recommendation models based on Mamba demonstrate superior advantages. TiSASRec (TiSASRec2020) enhances its performance by integrating time-aware capabilities via time difference encoding. Despite the theoretical FLOPS remaining in the order of $O(T^{2}N)$, the practical computational and memory demands exceed those of SASRec by more than four times. This substantial increase is attributed to the inherent computational constraints associated with time difference encoding. Consequently, while TiSASRec demonstrates improved time sensitivity, it necessitates a significantly higher computational overhead and memory allocation compared to SASRec, posing potential challenges in scalability and efficiency. In contrast, TiM4Rec preserves the benefits of the SSD architecture by employing a Time-aware Structured Masked Matrix characterized by linear computational complexity. This method preserves computational efficiency, requiring only $O(TN^{2})$FLOPs, and demands only an additional $O(T)$ level of FLOPs.
Notably, the generation of the time-aware masked matrix in TiM4Rec is solely based on scalar computations. This characteristic ensures that the additional computational complexity does not escalate in high-dimensional modeling scenarios, thereby preserving the superior matrix operation efficiency inherent to the SSD4Rec architecture. Our summarized results are shown in Table 1.
Table 1. Comparison of computational complexity between representative sequential recommendation methods, where ”M.M.” refers to ”Matrix Multiplication”, ”T.F” represents ”Time-aware additional FLOPs”
Architecture FLOPs T.F M.M.
SASRec Attention $T^{2}N$
LRURec LRU $\log(T)N^{2}$
Mamba4Rec SSM $TN^{2}$
SSD4Rec SSD $TN^{2}$
TiSASRec Attention $T^{2}N$ $T^{2}N$
TiM4Rec SSD $TN^{2}$ $T$
@@ 4.1.1. Datasets
  • MovieLens-1M (MovieLes2016): A dataset containing approximately 1 million user ratings for movies, collected from the MovieLens platform.
  • Amazon-Beauty (Amazon2014): The user review dataset collected in the beauty category on the Amazon platform was compiled up to the year 2014.
  • Kuai Rand (KuaiRand2022): Acquired from the recommendation logs of the application Kuaishou, the dataset includes millions of interactions involving items that were randomly displayed.
For each user, we sort their interaction records according to the timestamps, thereby generating an interaction sequence for each individual. And then retain only those users and items that are associated with a minimum of five interaction records, in accordance with the methodology established in prior research (SASRec2018). The statistics of these datasets are shown in Table 2.
@@
4.1.2. Baseline
To demonstrate the superiority of our proposed method, we select a set of representative sequential recommendation models to serve as baselines. Notably, even though SSD4Rec (SSD4Rec2024) has not been released as open-source, we implemented a variant termed SSD4Rec*, adhering to the traditional sequential recommendation framework to validate the efficacy of our time-aware approach. This implementation is strictly based on the SSD architecture and does not incorporate features such as variable-length sequences and bidirectional SSD methods proposed in SSD4Rec. The exclusion of these features is intentional, as they are generally regarded as universal enhancement techniques, particularly in the context of handling variable-length sequences.
  • Caser (TangW2018): A classical sequential recommendation model based on CNN.
  • GRU4Rec (GRU4Rec2016): A RNN-based method constructed by Gated Recurrent Units (GRU).
  • SASRec (SASRec2018): The first exploration model of applying Transformer in the field of sequential recommendation.
  • BERT4Rec (BERT4Rec2019): A bidirectional attention sequential recommendation model following the BERT (BERT2019) model paradigm.
  • TiSASRec (TiSASRec2020): A time-aware enhanced sequential recommendation model based on SASRec: Considered as a paradigm in time-aware sequential recommendation research.
  • LRURec (LRURec2024): A sequential recommendation model based on Linear Recurrent Units (LRU) (LRU2023) with Parallel Scan (Parallelizing2018) optimization.
  • Mamba4Rec (Mamba4Rec2024): A pioneering model that explores the application of the Mamba architecture in the domain of sequential recommendation.
  • SSD4Rec (SSD4Rec2024): The pioneering sequential recommendation model leveraging the SSD architecture, exploits its inherent advantages over the SSM architecture for high-dimensional modeling, and integrates variable-length sequence training techniques to enhance model performance.
Table 3. Sequential recommendation performance on the ML-1M, Beauty, and KuaiRand datasets. Here, R@k refers to Recall@K, N@K denotes NDCG@K, and M@K represents MRR@K. The best and second-best results are indicated in bold and underlined, respectively. ”Improv. SSD” signifies the performance improvement of TiM4Rec compared to the SSD4Rec* results. ”Improv. Sec” signifies the performance improvement of TiM4Rec compared to the second-best results. The asterisk in SSD4Rec* denotes that this is our own replicated version. The model enclosed in boxes represents the use of time aware enhancement methods.
Model Type & Model $\rightarrow$ CNN RNN Transformer LRU SSM SSD Improv. SSD Improv. Sec
Dataset Metric Caser GRU4Rec SASRec BERT4Rec TiSASRec LRURec Mamba4Rec SSD4Rec* TiM4Rec
MovieLens-1M R@10 0.2156 0.2985 0.3060 0.2800 0.3147 0.3002 0.3253 0.3199 0.3310 3.47% 1.75%
R@20 0.3083 0.3937 0.4050 0.3853 0.4250 0.4035 0.4354 0.4240 0.4338 2.31% -0.30%
R@50 0.4515 0.5397 0.5455 0.5377 0.5608 0.5455 0.5707 0.5649 0.5770 2.14% 1.10%
N@10 0.1157 0.1700 0.1754 0.1584 0.1834 0.1727 0.1891 0.1841 0.1932 4.94% 2.17%
N@20 0.1391 0.1941 0.2005 0.1852 0.2106 0.1987 0.2169 0.2104 0.2194 4.28% 1.15%
N@50 0.1675 0.2232 0.2285 0.2154 0.2376 0.2268 0.2436 0.2384 0.2477 3.90% 1.68%
M@10 0.0853 0.1307 0.1357 0.1214 0.1424 0.1337 0.1474 0.1425 0.1512 6.11% 2.58%
M@20 0.0917 0.1373 0.1425 0.1288 0.1498 0.1407 0.1551 0.1498 0.1584 5.23% 2.13%
M@50 0.0963 0.1420 0.1471 0.1337 0.1542 0.1452 0.1593 0.1542 0.1629 5.64% 2.26%
Amazon Beauty R@10 0.0402 0.0563 0.0851 0.0352 0.0802 0.0840 0.0838 0.0806 0.0854 5.96% 0.35%
R@20 0.0618 0.0856 0.1194 0.0563 0.1147 0.1191 0.1185 0.1146 0.1204 5.06% 0.84%
R@50 0.1090 0.1376 0.1759 0.0999 0.1734 0.1799 0.1802 0.1710 0.1800 5.26% -0.11%
N@10 0.0199 0.0297 0.0425 0.0172 0.0405 0.0427 0.0435 0.0423 0.0446 5.44% 2.53%
N@20 0.0253 0.0371 0.0511 0.0225 0.0491 0.0522 0.0522 0.0509 0.0533 4.72% 2.11%
N@50 0.0346 0.0473 0.0623 0.0311 0.0605 0.0640 0.0644 0.0620 0.0651 5.00% 1.09%
M@10 0.0138 0.0217 0.0294 0.0117 0.0282 0.0300 0.0311 0.0306 0.0321 4.90% 3.22%
M@20 0.0153 0.0237 0.0317 0.0132 0.0306 0.0326 0.0335 0.0329 0.0345 4.86% 2.99%
M@50 0.0167 0.0253 0.0335 0.0145 0.0325 0.0345 0.0354 0.0347 0.0363 4.61% 2.54%
Kuai Rand R@10 0.0801 0.1020 0.1055 0.0938 0.1057 0.1036 0.1094 0.1055 0.1109 5.12% 1.37%
R@20 0.1344 0.1659 0.1704 0.1537 0.1710 0.1663 0.1768 0.1717 0.1774 3.32% 0.34%
R@50 0.2561 0.3017 0.3074 0.2873 0.3060 0.3078 0.3154 0.3088 0.3202 3.69% 1.52%
N@10 0.0395 0.0564 0.0584 0.0510 0.0590 0.0570 0.0608 0.0588 0.0611 3.91% 0.49%
N@20 0.0531 0.0724 0.0747 0.0660 0.0753 0.0727 0.0777 0.0754 0.0779 3.32% 0.26%
N@50 0.0770 0.0911 0.1016 0.0923 0.1019 0.1005 0.1050 0.1024 0.1060 3.52% 0.95%
M@10 0.0273 0.0428 0.0443 0.0382 0.0450 0.0431 0.0461 0.0449 0.0463 3.12% 0.43%
M@20 0.0310 0.0471 0.0487 0.0422 0.0494 0.0473 0.0508 0.0494 0.0508 2.83% 0.00%
M@50 0.0347 0.0513 0.0529 0.0464 0.0536 0.0517 0.0552 0.0536 0.0552 2.99% 0.00%
4.1.3. Evaluation Metrics